OpenStudy (anonymous):

Find the inverse Laplace transform of F(s)=1/(s^4(s^2+1)) by using the convolution theorem. (I got L^-1 (1/s^4)*L^-1 (1/(s^2+1)) but how do I find the answer?)

OpenStudy (anonymous):

@robtobey

OpenStudy (loser66):

\[L^{-1}\frac{1}{s^4}= \frac{t^3}{6}\] \[L^{-1}\frac{1}{s^2+1}= sin t\] convolution theorem said that \[f(t) =\int_0^t \frac {(t-taute)^3}{6}sin(taute)d(taute)\] then take integral of this. That's it, stubborn girl. XD

OpenStudy (anonymous):

How did you find t^3/6?

OpenStudy (loser66):

laplace transform table.

OpenStudy (anonymous):

Wait a minute. Let me check the table.

OpenStudy (loser66):

formula #4, stubborn girl

OpenStudy (anonymous):

Don't call me stubborn, you mean woman.

OpenStudy (loser66):

so what??? you call me "mean woman" all the time, Why don't I feel bad? Hey, people study math is not that weak, girl. Ok, when you are beaten, wait for a good chance to beat back, not upset, girl.

OpenStudy (anonymous):

hahahaha, mean woman...

OpenStudy (loser66):

ok, I like a strong one. fair enough.